### v2

# FG-PINNs: frequency-guided physics-informed neural networks for solving PDEs with high frequency components

Abstract.
In this work, we propose the frequency-guided physics-informed neural networks (FG-PINNs), specifically designed for solving partial differential equations (PDEs) with high-frequency components. The core of the proposed algorithm lies in utilizing high-frequency information obtained from PDEs to guide the neural network in rapidly approximating the high-frequency components of the solution. The FG-PINNs consist of two subnetworks, including a high-frequency subnetwork for capturing high-frequency components and a low-frequency subnetwork for capturing low-frequency components. The key innovation in the high-frequency subnetworks is to embed prior knowledge for high-frequency components into the network structure. For nonhomogeneous PDEs ($f(x)\neq c,c\in R$), we embed the source term with high-frequency components into the neural network. For homogeneous PDEs, we embed the initial/boundary conditions with high-frequency components into the neural network. Based on spectral bias, we use a fully connected neural network as the low-frequency subnetwork to capture low-frequency components of the solution. A series of numerical examples demonstrate the effectiveness of the FG-PINNs, including the one-dimensional heat equation (relative $L^{2}$ error: $O(10^{-4})$), the nonlinear wave equations (relative $L^{2}$ error: $O(10^{-4})$) and the two-dimensional heat equation (relative $L^{2}$ error: $O(10^{-3})$).
Key words and phrases:
Physics-informed neural networks; Spectral bias; Nonhomogeneous PDEs
Corresponding author.
1. Introduction
In scientific computing, deep neural networks (DNNs) are widely used to solve various PDEs [1, 2, 3, 4, 5, 6, 7]. Although many important results [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have been achieved in solving PDEs by DNNs, there are still many problems that need to be addressed, such as spectral bias (frequency principle) [19]. This bias is manifested in the tendency of DNNs to prioritize learning low-frequency components over high-frequency ones when fitting functions or solving PDEs. The existence of spectral bias has been experimentally demonstrated in many literatures [20, 21, 22], and Tancik et al. use the neural tangent kernel (NTK) [23] to provide corresponding theoretical support. Inspired by insights from NTK, multi-scale random Fourier feature mapping is adopted in physics-informed neural networks (PINNs) to enhance robustness and accuracy [24, 25]. Additionally, PINNs utilize derivatives of the solution to define the loss function, which may accelerate the convergence of high-frequency components during training. Examples provided by Lu et al. [21] demonstrate that differentiation can accelerate the convergence of high-frequency components.
In the field of computer vision, numerous studies have fully leveraged the learning disparities of deep neural networks at different frequencies, such as in classification [32], image recovery [33], and super-resolution [34]. By embedding frequency information into neural network architectures, the training speed and generalization of models have been significantly improved. To address the spectral bias, Cai et al. [26] proposed the phase shift DNNs (PhaseDNNs). By employing phase shift techniques in the frequency domain, PhaseDNNs convert high-frequency data into low-frequency data. Once the training is completed, the data is transformed back to the original frequency through an inverse phase shift. However, as the number of Fourier terms increases exponentially with dimensions, PhaseDNNs face difficulties in dealing with high-dimensional problems. As another way to overcome spectral bias, multiscale DNNs [27, 28] use radial scaling techniques to transform the frequency of the data, as well as an activation function with local frequencies to learn the frequency components of the objective function. MscaleDNNs and PhaseDNNs consist of a series of subnetworks, each of which is responsible for approximating a specific range of frequencies. Huang et al. [29] proposed a frequency-adaptive multi-scale DNNs (frequency-adaptive MscaleDNNs) to reduce the parameter dependency of MscaleDNNs. The frequency-adaptive MscaleDNNs employ a multi-stage training strategy. In the first stage, MscaleDNNs are used to obtain a pre-trained solution. The frequency coefficient distribution of the pre-trained solution is acquired via Fourier transform, and these corresponding frequencies are then utilized to construct a new neural network with multiple subnetworks. This new network is trained to achieve a more accurate approximation, iterating this process until training is complete.
When solving PDEs, deep neural networks are constrained by inherent spectral bias, struggling to accurately approximate the solution with a wide range of frequencies, particularly in capturing high-frequency components. We observe that source terms and initial/boundary conditions usually provide information about high-frequency components. To fully utilize this information, we developed frequency-guided physics-informed neural networks (FG-PINNs) for solving PDEs with high-frequency components. The key innovations in FG-PINNs are summarized as follows:
Dual Network Structure: FG-PINNs consist of two sub-networks, including a high-frequency sub-network and a low-frequency sub-network. The high-frequency sub-network is responsible for approximating high-frequency components, while the low-frequency sub-network handles the approximation of low-frequency components.
Frequency Prior: The source term and initial/boundary value conditions of the problem to be solved usually provide prior knowledge about the high-frequency components. By embedding this prior knowledge about high-frequency components into the neural network, the neural network can rapidly approximate the high-frequency components of the solution.
The rest of the paper is organized as follows: In section 2, we briefly review the PINNs for solving PDEs. In section 3, we explain the network structure of FG-PINNs in detail. In section 4, the different PDEs with high-frequency components are used to test the approximation ability of FG-PINNs. Finally, the advantages and inadequacies of FG-PINNs are summarized in section 5.
2. Review of physics-informed neural networks
In this section, we briefly review the PINNs for solving PDEs. Considering the following problem:
[…]
where $\Omega\in\mathbb{R}^{N}$ denotes the computational domain, and $\Gamma$ represents a differential operator containing temporal and spatial derivatives. The corresponding residual functions of problem (2.1) are defined as:
[…]
where $(\textbf{x},t)$, $(\textbf{x}_{b},t_{b})$ and $(\textbf{x}_{0},0)$ denote the interior point, the boundary point, and the initial point, respectively. $u_{nn}(\textbf{x},t,\theta)$ denotes a fully connected neural network with $M$ hidden layers defined as follows:
$\Phi^{0}(\bar{\textbf{x}})=\bar{\textbf{x}},$ (2.3)
$\Phi^{m}(\bar{\textbf{x}})=\sigma(W^{m}\Phi^{m-1}(\bar{\textbf{x}})+b^{m}),\qquad 1\leq m\leq M,$
$u_{nn}(\bar{\textbf{x}},W)=W^{M+1}\Phi^{M}(\bar{\textbf{x}})+b^{M+1},$
where $\bar{\textbf{x}}=\{\textbf{x},t\}$ and $\Phi^{m}(\bar{\textbf{x}})$ represent the input and hidden state of the neural network, respectively. $\sigma(\cdot)$ is a nonlinear activation function. $\theta=\{W^{m},b^{m}\}_{m=1}^{M+1}$ denote all the trainable parameters. Then, a neural network approximation for problem (2.1) is obtained by minimizing the following loss function:
[…]
where $\{\tau,\tau_{b},\tau_{0}\}$ denote the weights of each of the components in the loss function.
3. Frequency-guided physics-informed neural networks
The spectral bias inherent in deep neural networks makes it difficult for them to accurately approximate solutions with a wide range of frequencies. This bias leads the networks to preferentially learn the low-frequency components of the solution, while the convergence of high-frequency components is extremely slow. It is worth noting that for PDEs with high-frequency components, their source terms, initial/boundary conditions usually contain prior information about the high-frequency components in the solution. However, PINNs and their variants fail to fully utilize this valuable frequency information. To fully utilize this prior information on frequency components, we propose the frequency-guided physics-informed neural networks (FG-PINNs) for solving PDEs with high-frequency components. The core idea in FG-PINNs is to extract and integrate prior frequency information from the source terms and initial/boundary conditions to guide the network in rapidly learning the high-frequency components of the solution. The FG-PINNs consist of two subnetworks, including a high-frequency subnetwork for capturing high-frequency components and a low-frequency subnetwork for capturing low-frequency components. The forward propagation rule for high-frequency subnetwork with $N^{*}$ module is as follows:
$H^{0}=\textbf{x},$ (3.5)
$Q^{n}=W^{n}_{Q}H^{n-1},\;K^{n}=W^{n}_{K}H^{n-1},\;V^{n}=W^{n}_{V}H^{n-1},$
$A=\sigma({Q^{n}\odot K^{n}})\odot V^{n},$
$H=(1-\alpha^{n})A+\alpha^{n}\hbar(\textbf{x}),$
$H^{n}=\sigma(W^{n}H+b^{n}),\quad 1\leq n\leq N^{*},$
$u_{nn}^{h}(\textbf{x})=W^{N^{*}+1}H^{N^{*}}+b^{N^{*}+1},$
where $H^{n}$ denotes the output of the n-th module. $\odot$ is the element-wise multiplication. $\{\textbf{x},\;u_{nn}^{h}(\textbf{x})\}$ represent the input and output of the high-frequency subnetworks, respectively. $Q^{n},K^{n}$ and $V^{n}$ are employed to learn different features and are integrated within the hidden layer $H$. $\hbar(\textbf{x})$ represents the source items or initial/boundary value conditions. In hidden layer $H$, a weighting factor $\alpha^{n}$ is needed to control the contribution weight of $\hbar(\textbf{x})$. In nonhomogeneous problems with high frequency components, the value of the source term is usually very large. This may lead to problems such as gradient vanishing. Therefore, the source term needs to be normalized. In this work, we used an improved maximum normalization:
$\hbar(\textbf{x})=\frac{f(\textbf{x})}{K^{*}|f(\textbf{x})|_{max}},\quad\textbf{x}\in\Omega,$ (3.6)
where $K^{*}\geq 1$ is a manually adjustable hyperparameter.
The high-frequency subnetwork is employed to handle the high-frequency components of the solution. Inspired by spectral bias, we use a fully connected neural network to handle the low-frequency components, i.e., the low-frequency subnetwork. Finally, the high-frequency subnetwork and low-frequency subnetwork are integrated to generate the final output:
$u_{nn}^{*}(\textbf{x},\theta)=u_{nn}^{h}(\textbf{x},\theta)+u_{nn}^{l}(\textbf{x},\theta),$ (3.7)
where $u_{nn}^{h}(\cdot)$ and $u_{nn}^{l}(\cdot)$ represent the high-frequency subnetwork and low-frequency subnetwork, respectively. The network structure of FG-PINNs is shown in Figure 1. Then the loss function $L(\textbf{x},t,\theta)$ of problem (2.1) is reformulated as:
$L^{*}(\textbf{x},t,\theta)$ $=\tau\frac{1}{N}\sum_{i=1}^{N}r^{*}(\textbf{x}^{i},t^{i},\theta)+\tau_{b}\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}r^{*}_{b}(\textbf{x}_{b}^{i},t_{b}^{i},\theta)+\tau_{0}\frac{1}{N_{0}}\sum_{i=1}^{N_{0}}r^{*}_{0}(\textbf{x}^{i}_{0},0,\theta),$ (3.8)
$=\tau\frac{1}{N}\sum_{i=1}^{N}|\Gamma u_{nn}^{*}(\textbf{x}^{i},t^{i},\theta)-f(\textbf{x}^{i},t^{i})|^{2}$
$+\tau_{b}\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}|u_{nn}^{*}(\textbf{x}^{i}_{b},t^{i}_{b},\theta)-h(\textbf{x}^{i}_{b},t^{i}_{b})|^{2}$
$+\tau_{0}\frac{1}{N_{0}}\sum_{i=1}^{N_{0}}|u_{nn}^{*}(\textbf{x}^{i}_{0},0,\theta)-g(\textbf{x}^{i}_{0})|^{2}.$
By minimizing $L^{*}(\textbf{x},t,\theta)$, we obtain an approximation $u_{nn}^{*}(\textbf{x},t,\theta)$ of problem (2.1). In order to obtain the optimal parameters $\theta^{*}=\{W^{*},b^{*}\}$, we use the gradient descent method to update the parameters. The update rules for step $k$ are as follows:
$W^{k+1}=$ $W^{k}-\eta^{k}\frac{\partial L^{*}(\textbf{x},t,W^{k},b^{k})}{\partial W^{k}},$ (3.9)
$b^{k+1}=$ $b^{k}-\eta^{k}\frac{\partial L^{*}(\textbf{x},t,W^{k},b^{k})}{\partial b^{k}},$
where $\eta^{k}$ denotes the step size of the $k$-th iteration.
Figure 1. The neural network structure of FG-PINNs.
The solution containing high-frequency components may generate a large-magnitude source term. This leads to a significant scale disparity between the residual component $r^{*}(\textbf{x},t,\theta)$ and other loss components (e.g., initial/boundary conditions) in the loss function. To mitigate this imbalance, we propose a new weighting method:
$\tau=e^{-K_{\tau}\mid f(\textbf{x})/f(\textbf{x})_{max}\mid},$ (3.10)
where $K_{\tau}$ is a manually adjustable hyperparameter, usually set to $1$. This weighting method assigns small weight to the residual with large source term value, to balance the difference between $r^{*}(\textbf{x},t,\theta)$ and the other loss function components. Finally, we summarize the core differences between the FG-PINNs and the existing algorithms:
Differences in Network Architecture: PhaseDNNs, MscaleDNNs, and Frequency-adaptive MscaleDNNs rely on parallel multi-subnetwork structures. In contrast, FG-PINNs adopt a dual-network architecture, where low-frequency and high-frequency subnetworks collaborate to capture the frequency components of the solution.
Independence from Scale Parameters: PhaseDNNs and MscaleDNNs require manual presetting of critical scale parameters, a process that is time-consuming and experience-dependent. Notably, FG-PINNs completely eliminate the need for manual scale parameter tuning.
Significant Improvement in Convergence Efficiency: PhaseDNNs and MscaleDNNs typically require tens of thousands to hundreds of thousands of iterations, while Frequency-adaptive MscaleDNNs can require hundreds of thousands due to their multi-stage training. In sharp contrast, FG-PINNs generally converge within just thousands of iterations, demonstrating a substantial efficiency advantage.
However, we should note that FG-PINNs have the following limitation:
The limitation for FG-PINNs: When $f\equiv c$ $(c\in R)$ and the initial/boundary conditions do not provide information about the high-frequency components, it will be difficult for FG-PINNs to provide an accurate approximation for PDEs with high-frequency components.
4. experiments
In this section, we evaluate the performance of FG-PINNs in solving PDEs with high-frequency components. In all simulations, we use Adam [30] to optimize the neural network, where the learning rate is set to 0.001. the parameters $K^{*}=2,\;K_{\tau}=1$. In particular, we set the global random number seed to 1234. The relative $L^{2}$ error is used to evaluate the performance of the proposed model:
$E_{L}=\frac{\sqrt{\sum_{i=1}^{N}|u(\textbf{x}_{i})-u_{nn}^{*}(\textbf{x}_{i},\theta)|^{2}}}{\sqrt{\sum_{i=1}^{N}|u(\textbf{x}_{i})|^{2}}},$ (4.11)
where $u(\textbf{x}_{i})$ is the reference solution.
Example 1.
We solve the one-dimensional Poisson equation using FG-PINNs:
[…]
where $\Omega=[0,2\pi]$, and the exact solution:
$u(x)=(sin(x)+sin(10x)+sin(100x))/3.$ (4.13)
In this test, we use 2,000 interior points sampled by the Latin hypercube sampling (LHS) [31] and two boundary points as training points. A fully connected neural network with 4 hidden layers serves as the low-frequency subnetwork. We use 4 modules with 48 neurons to construct the high-frequency subnetwork. The weight $\tau_{b}$ is set to 100. The maximum number of iterations is set to 5,000.
Figure 2. Example 1: Poisson equation (4.12). A: Exact solution and predicted solution. B: Absolute error. C: Exact spectrum and predicted spectrum. D: Convergence of the amplitude for each frequency during the training process.
Figure 2 (A) shows the exact solution and the predicted solution obtained by FG-PINNs. It can be observed that FG-PINNs provide an accurate approximation. The absolute error is displayed in Figure 2 (B), where the maximum absolute error is $3.56\times 10^{-3}$. Figure 2 (C) shows the exact spectrum and the predicted spectrum, demonstrating that FG-PINNs accurately capture the frequency components of the exact solution. Figure 2 (D) displays the convergence evolution of the amplitude for each frequency during training. Numerical results indicate that FG-PINNs rapidly converge to the amplitude of each frequency. Notably, the most rapid convergence is achieved for the high-frequency component $sin(100x)$ using FG-PINNs, followed by $sin(10x)$. Figure 3 shows the spectrum of the two subnetworks. The numerical results show that the high-frequency subnetwork accurately captures the high-frequency components, and the low-frequency sub-network accurately captures the low-frequency components.
Figure 3. Example 1: Poisson equation (4.12). A: $u_{nn}^{h}(x)$ spectrum. B: $u_{nn}^{l}(x)$ spectrum.
Example 2.
We consider the following one-dimensional heat equation:
[…]
The exact solution:
$u(x,t)=e^{-t}sin(500\pi x).$ (4.15)
In this test, problem (4.14) is a homogeneous PDE. We note that the initial condition provides information about the high-frequency component. Therefore, the initial condition is embedded into the high-frequency subnetwork. 5,000 interior points and 2,000 boundary points are used for network training. We use a fully connected neural network with 4 hidden layers as the low-frequency subnetwork, with 48 neurons in each hidden layer. Three modules with 48 neurons are used to construct the high-frequency subnetwork. For each subnetwork, the $sin(x)$ function served as the activation function. The maximum number of iterations is set to 3,000.
Figure 4 (A) and (B) show the exact solution and the predicted solution, respectively. The numerical results show that FG-PINNs provide an accurate approximation. Figure 4 (C) displays the absolute error of FG-PINNs, where the maximum absolute error reaches $O(10^{-3})$. Figure 4 (D) displays the relative $L^{2}$ error curve during the training process. This curve reveals a key observation: during the initial stages of training, the relative $L^{2}$ error exhibits an extremely rapid decline. This phenomenon indicates that FG-PINNs can rapidly approximate the frequency components of the solution.
We test the performance of the model at different numbers of modules and training points. Table 1 provides a detailed comparison. The numerical results in Table 1 leads to a clear conclusion: As the number of modules used to build the high-frequency subnetwork increases and the number of sample points used for training grows, the accuracy of the approximated solution obtained by FG-PINNs continues to improve significantly. This fully demonstrates that performance of FG-PINNs can be effectively improved by increasing the complexity of the model and the number of training points.
Table 1. Relative $L^{2}$ errors at different numbers of modules and training points in Example 2
    Interior points Modules    1    2    3    4
   1000    $5.42\times 10^{0}$    $9.12\times 10^{-3}$    $6.10\times 10^{-3}$    $7.94\times 10^{-4}$
   3000    $1.81\times 10^{0}$    $1.36\times 10^{-3}$    $5.07\times 10^{-4}$    $4.21\times 10^{-4}$
   5000    $3.27\times 10^{-2}$    $8.36\times 10^{-4}$    $3.82\times 10^{-4}$    $4.16\times 10^{-4}$
Figure 4. Example 2: Heat equation (4.14). A: Exact solution. B: Predicte solution. C: Absolute error. D: Relative $L^{2}$ error.
Table 2. Relative $L^{2}$ errors for different activation functions in Example 2
   Activation functions    sin(x)    cos(x)    tanh(x)    sigmoid(x)
   Relatvie $L^{2}$ errors    $3.82\times 10^{-4}$    $6.58\times 10^{-4}$    $4.29\times 10^{-4}$    $7.95\times 10^{-4}$
Example 3.
We consider the following nonhomogeneous wave equation:
$\begin{split}&u_{tt}(x,y,t)-(u_{xx}(x,y,t)+u_{yy}(x,y,t))+u(x,y,t)^{3}=f(x,y,t),\,\quad[x,y,t]\in\Omega\times[0,T],\\ &u(x,y,t)=g(x,y,t),\,\;\quad\quad\quad\quad\quad\;\;\,\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad[x,y]\in\partial\Omega,t\in[0,T],\\ &u(x,y,0)=h(x,y),\,\,\quad\quad\quad\quad\quad\;\,\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad[x,y]\in\Omega,\\ &u_{t}(x,y,0)=m(x,y),\,\;\;\;\;\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad[x,y]\in\Omega,\end{split}$ (4.16)
where $\Omega=[0,1]^{2}$ and $T=1$, the exact solution:
$u(x,y,t)=cos(50\pi x)cos(100\pi y)sin(t).$ (4.17)
In this test, we employed LHS to generate 3,500 training points, comprising 2,000 interior points, 1,200 boundary points, and 300 initial points. The weights are set to $t_{b}$ = 100 and $t_{0}$ = 100. We utilized a fully connected neural network with 4 hidden layers as the low-frequency subnetwork, each hidden layer containing 48 neurons. Three modules, each with 48 neurons, are used to construct the high-frequency subnetwork. For each subnetwork, the $cos(x)$ function served as the activation function, and the maximum number of iterations is set to 1,000.
Figure 5 (A) and Figure 5 (B) show the exact solution (t=0.5) and the predicted solution (t=0.5), respectively. Figure 5 (C) displays the absolute error of FG-PINNs, where the maximum absolute error reaches $8.06\times 10^{-4}$. The numerical results demonstrate that FG-PINNs provides an accurate approximation. Figure 5 (D) shows the relative $L^{2}$ error curve during the training process. After approximately 200 iterations, the relative $L^{2}$ error of FG-PINNs decreases to $O(10^{-3})$. This indicates that FG-PINNs can rapidly capture both the high-frequency and low-frequency components of the solution.
To systematically evaluate the sensitivity of the FG-PINNs model performance to different activation functions, we conducted a series of controlled experiments. Table 3 provides a detailed comparison. It can be observed that $cos(x)$ achieved the best approximation in this case. We tested the model’s performance under varying numbers of modules and training points. Table 4 shows detailed numerical results. As the number of modules used to build the high-frequency subnetwork increases and the number of sample points used for training grows, the accuracy of the approximation obtained by FG-PINNs is significantly improved. A series of numerical experiments verifies the effectiveness of the proposed method.
Figure 5. Example 3: Wave equation (4.16). A: Exact solution (t=0.5). B: Predicte solution (t=0.5). C: Absolute error. D: Relative $L^{2}$ error.
Table 3. Relative $L^{2}$ errors for different activation functions in Example 3
   Activation functions    sin(x)    cos(x)    tanh(x)    sigmoid(x)
   Relatvie $L^{2}$ errors    $6.404\times 10^{-3}$    $1.01\times 10^{-3}$    $9.27\times 10^{-3}$    $8.95\times 10^{-3}$
Table 4. Relative $L^{2}$ errors at different numbers of modules and training points in Example 3
    Interior points Modules    1    2    3    4
   1000    $1.43\times 10^{-1}$    $1.57\times 10^{-2}$    $9.58\times 10^{-3}$    $6.45\times 10^{-3}$
   3000    $5.28\times 10^{-2}$    $6.75\times 10^{-3}$    $1.01\times 10^{-3}$    $2.98\times 10^{-3}$
   5000    $54\times 10^{-2}    4.4$$41\times 10^{-3}$    $2.27\times 10^{-3}$    $3.12\times 10^{-3}$
Figure 6. Example 4: Wave equation (4.18). A: Exact solution (t=0.5). B: Predicte solution (t=0.5). C: Absolute error. D: Relative $L^{2}$ error.
Table 5. Relative $L^{2}$ errors at different numbers of modules and training points in Example 4
    Interior points Modules    1    2    3    4
   1000    $7.96\times 10^{-2}$    $9.34\times 10^{-3}$    $1.78\times 10^{-3}$    $1.71\times 10^{-3}$
   2000    $1.48\times 10^{-2}$    $5.41\times 10^{-3}$    $1.77\times 10^{-3}$    $1.32\times 10^{-3}$
   3000    $7.41\times 10^{-3}$    $3.12\times 10^{-3}$    $1.08\times 10^{-3}$    $9.41\times 10^{-4}$
Example 4.
We consider the following non-homogeneous wave equation:
$\begin{split}&u_{tt}(x,y,t)-(u_{xx}(x,y,t)+u_{yy}(x,y,t))+u_{t}(x,y,t)^{2}=f(x,y,t),\quad[x,y,t]\in\Omega\times[0,T],\\ &u(x,y,t)=g(x,y,t),\;\;\quad\quad\quad\quad\quad\;\;\;\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad[x,y]\in\partial\Omega,t\in[0,T],\\ &u(x,y,0)=h(x,y),\,\,\,\quad\quad\quad\quad\quad\;\,\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad[x,y]\in\Omega,\\ &u_{t}(x,y,0)=m(x,y),\,\;\;\;\;\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad[x,y]\in\Omega,\end{split}$ (4.18)
where $\Omega=[0,1]^{2}$ and $T=1$, the exact solution:
$u(x,y,t)=e^{-2t}sin(\pi x)sin(100\pi y).$ (4.19)
In this test, the training points and the weights within the loss function are consistent with those used in Example 3. We use a fully connected neural network with 4 hidden layers as the low-frequency subnetwork, each layer containing 48 neurons. Four modules, each with 48 neurons, are used to construct the high-frequency subnetwork. The $sin(x)$ function served as the activation function for each subnetwork and the maximum number of iterations is set to 6,000.
Figure 6 (A) and (B) display the exact solution (t=0.5) and the predicted solution (t=0.5), respectively. It can be found that FG-PINNs provide an accurate approximation. The absolute error is presented in Figure 6 (C), where the maximum absolute error is $6.31\times 10^{-4}$. The relative $L^{2}$ error is shown in Figure 6 (D). Similarly, during the early stages of training, the relative $L^{2}$ error decreases rapidly to the order of $O(10^{-3})$. After 6000 iterations, FG-PINNs yield an approximation with relative $L^{2}$ error of $O(10^{-4})$. Table 6 provides a detailed performance comparison of different activation functions under the same experimental configuration. The numerical results indicate that $sigmoid(x)$ is the most suitable and optimal activation function for this case. Furthermore, Table 5 provides a detailed comparison concerning different numbers of modules and training points. As the number of modules and training points increases, the accuracy of the approximate solution obtained by the FG-PINNs is significantly improved.
Table 6. Relative $L^{2}$ errors for different activation functions in Example 4
   Activation functions    sin(x)    cos(x)    tanh(x)    sigmoid(x)
   Relatvie $L^{2}$ errors    $9.41\times 10^{-4}$    $2.94\times 10^{-3}$    $1.15\times 10^{-3}$    $5.99\times 10^{-4}$
Example 5.
Finally, consider a nonhomogeneous anisotropic two-dimensional heat equation:
$u_{t}(x,y,t)=(k_{11}u_{xx}+k_{12}u_{xy}+k_{21}u_{yx}+k_{22}u_{yy})(x,y,t)+f(x,y,t),[x,y,t]\in[0,1]^{3},$ (4.20)
where $k_{ij}\ (i,j=1,2)$ are the thermal conductivity coefficients in the anisotropic material. $K=(k_{ij})_{2\times 2}$ denotes the thermal conductivity matrix, satisfying the Onsager reciprocal relation $k_{12}=k_{21}$ and $\det(K)>0$. The exact solution is:
$u(x,y,t)=e^{-2t}(cos(x+50\pi y)+cos(-3x+4y)).$ (4.21)
Temperature distribution at a given moment (t=1):
$u(x,y,1)=e^{-2}(cos(x+50\pi y)+cos(-3x+4y)).$ (4.22)
The initial temperature distribution is unknown, and we need to use the temperature distribution at $t=1$ to reconstruct the initial temperature, i.e., a time-reversed problem. Note that time-reversed problems are typically ill-posed.
In this case, we set the thermal conductivity coefficients $k_{11}=2,k_{22}=2.5,k_{12}=k_{21}=1$. We employ three modules with 64 neurons each to construct the high-frequency subnetwork. A fully connected neural network with four hidden layers, each containing 64 neurons, serves as the low-frequency subnetwork. We generated 4500 training points using LHS, comprising 3,000 interior points, 1,200 boundary points, and 300 final time points (t=1). The maximum number of iterations was set to 6,000.
Figure 7 (B) and (C) depict the temperature distribution and prediction at the initial time, respectively. The numerical results show that FG-PINNs provide an accurate prediction. Figure 7 (A) shows the corresponding absolute error, where the maximum absolute error and the relative $L^{2}$ error are $1.77\times 10^{-2}$ and $6.21\times 10^{-3}$, respectively. The numerical results indicate that FG-PINNs effectively reconstruct the initial temperature distribution. Figure 8 (A) and (B) display the exact solution (t=0.5) and the prediction (t=0.5) of FG-PINNs, respectively. The absolute error is shown in Figure 8 (C), with maximum absolute error and relative $L^{2}$ error of $2.98\times 10^{-3}$ and $1.53\times 10^{-3}$, respectively. Figure 8 (D) illustrates the relative $L^{2}$ error curve. It can be observed that the relative $L^{2}$ error curve oscillates significantly during the training process, which may be attributed to the ill-conditioned nature of the problem itself. After 6,000 iterations, FG-PINNs yield an accurate approximation.
Under the same experimental configuration, we evaluated the performance of different activation functions. All numerical results are summarized in Table 8, where $sin(x)$ achieves the best approximation. Finally, we tested the performance of FG-PINNs with varying numbers of modules and training points. Table 7 provides detailed numerical results. As the number of modules and training points increases, FG-PINNs provide a more accurate approximation.
Table 7. Relative $L^{2}$ errors at different numbers of modules and training points in Example 5
    Interior points Modules    1    2    3    4
   1000    $2.58\times 10^{-1}$    $1.31\times 10^{-2}$    $8.45\times 10^{-3}$    $4.09\times 10^{-3}$
   2000    $3.76\times 10^{-2}$    $7.33\times 10^{-3}$    $4.04\times 10^{-3}$    $2.83\times 10^{-3}$
   3000    $8.07\times 10^{-3}$    $3.18\times 10^{-3}$    $1.53\times 10^{-3}$    $1.59\times 10^{-3}$
Table 8. Relative $L^{2}$ errors for different activation functions in Example 5
   Activation functions    sin(x)    cos(x)    tanh(x)    sigmoid(x)
   Relatvie $L^{2}$ errors    $1.53\times 10^{-3}$    $2.13\times 10^{-3}$    $2.95\times 10^{-3}$    $6.13\times 10^{-3}$
Figure 7. Example 5: Heat equation (4.20). A: Absolute error. B: Initial temperature distribution. C: Prediction.
Figure 8. Example 5: Heat equation (4.20). A: Exact solution (t=0.5). B: Predicte solution (t=0.5). C: Absolute error. D: Relative $L^{2}$ error.
5. Summary and discussion
In this work, we developed frequency-guided physics-informed neural networks (FG-PINNs) for solving PDEs with high-frequency components. The proposed algorithm relies on prior knowledge about high-frequency components obtained from PDEs. It aims to use this prior knowledge to guide neural networks in rapidly approximating the solution’s high-frequency components. FG-PINNs consist of two subnetworks: a high-frequency subnetwork and a low-frequency subnetwork. We employ the high-frequency subnetwork, incorporating the source terms or initial/boundary conditions with high-frequency components, to capture the high-frequency components of the solution. Inspired by spectral bias, we utilize a fully connected neural network to capture the low-frequency components of the solution. To evaluate the performance of FG-PINNs, we conducted a series of tests. The numerical results demonstrate that FG-PINNs can provide an accurate approximation for PDEs with high-frequency components.
Compared to existing methods that require multiple subnetworks, FG-PINNs only need two subnetworks, thereby saving the associated computational overhead. Meanwhile, FG-PINNs exhibit rapid convergence, as evidenced by the aforementioned series of experiments. This accelerated convergence stems from the fact that FG-PINNs incorporate the high-frequency information of the solution. Therefore, both the high-frequency and low-frequency components of the solution achieve rapid convergence when using FG-PINNs to solve PDEs with high-frequency components. In contrast, existing methods typically require tens of thousands to hundreds of thousands of iterations to obtain an accurate approximation. However, it should be noted that FG-PINNs are extremely dependent on prior information about high-frequency components obtained from PDEs.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This work was supported by NSFC Project (12431014), Project of Scientific Research Fund of the Hunan Provincial Science and Technology Department (2024ZL5017) and Guizhou Provincial Science and Technology Projects, China (No. QKHJC-ZK[2023]YB036).